when we placed a boy 4 feet from the fulcrum, and a girl 12 feet from the fulcrum on the side of the teeter board, they balance a man weighing 150 pounds who is seated 6 feet from the fulcrum on the other side. also, the board is balanced if the boy moves to 8 feet and the girl to 6 feet from the fulcrum on their side. Find their weights
Answer by mananth(16946)

let weight of boy be x --- 4 feet from fulcrum
weight of girl be y ---12 feet from fulcrum
balances 150*6
4x+12y=150*6..........1

Similarly
8x+6y= 150*6.........2
multiply (1) by -2
we get
-8x-24y=-1800
add this to (2)
-18y=-900
y=50 pounds weight of girl
plug value of y in (2)
8x+300=900
8x=600
x=75 pounds weight of boy
m.ananth@hotmail.ca